# PQMass: Probabilistic Assessment of the Quality of Generative Models using Probability Mass Estimation

### v2

@@ 4 Limitations
PQMass was shown to perform well with large number of samples. The statistics, however, will be too noisy for practical use if this number is too limited. Similarly, the number of regions is important: while our ablation study (§ B) shows that the results are robust to various choices, in the limits $n_{R}=1$ and $n_{R}\rightarrow N$ (where $N$ is the number of samples), sensitivity will be limited.
@@
Finally, PQMass is only as good as the distance metric used, as shown in § G. For some modalities (for example, natural language), it is nontrivial to select a meaningful distance metric between samples.
@@ 5 Conclusion
We have shown the performance of PQMass on a variety of synthetic tasks, as well as on comparing sampling methods, comparing generative models of images, and detecting hidden signals in time series. Given the versatility (further shown in the additional experiments in § F) and low computational cost of PQMass, it can serve as a valuable tool for evaluating the quality and performance of generative models and sampling methods.
@@ Appendix A Null tests
Figure 9: Null tests, similar to the one shown in Fig. 2, but for two different distributions: Samples from MNIST (left) and a VAE trained on MNIST (right). The main result of our paper holds for samples from any distribution as long as we are comparing two sets of independent samples from the same distribution. Note that, on MNIST, the fit is not perfect due to the limited number of data points.
The main result of this paper, shown in § 2, was shown in practice in § 3.1 for a high-dimensional mixture of Gaussians. In this section, we show that it works in even more complex and high-dimensional distributions. First, we test it on the MNIST data: Using the MNIST observations themselves as samples from some underlying probability distribution. We put together the MNIST train and test data, leading to $70,000$ images. We then split them into two subsets of $35,000$ images each. To repeat the null test of § 3.1, we need independent samples at every iteration (note that this is different from the error bars reported in the rest of the experiments, which just arise from repeating with random tessellations, but always the same samples). Therefore, for each iteration, we take $1,000$ images from each dataset, and perform the PQM test with $n_{R}=100$, leading to 35 data points. We plot their histogram in the left panel of Fig. 9. While noisy, due to the very few data points, we see that the histogram does follow a chi-squared distribution with $n_{R}-1$ degrees of freedom, as expected.
For our second experiment, we use one of the MNIST generative models described in § 3.3. In this case, we can generate as many samples as we want. Because we are comparing the generative model to itself, the similarity of the generated samples to the MNIST train or test data is irrelevant for this particular test.
In practical applications, the number of samples available from real data might be limited, and in such cases, the PQMass distribution can be approximated by reusing the samples and only retessellating, which also yields good results. In practice, we have observed that, for typical sample sizes, the match between the $\chi^{2}_{PQM}$ distribution obtained by retessellating without resampling and the theoretical $\chi^{2}$ distribution is very close and more than adequate for testing the null hypothesis (we have indicated in the captions of the figures and the tables the specific experiments where this was done). However, in the cases where reusing the samples between retessellation, the retessellations are not independent since the samples are shared between each test, which might affect the obtained distribution. We explore further the impact of this approximation in § I.
@@ Appendix B Ablation study
Table 3: Ablation study for CIFAR-10. We repeat the experiment of § 3.3, varying the number of reference points, as well as changing the distance metric to L1. We show the results for the frequentist version of PQMass, sorted in order of decreasing $\chi^{2}_{\rm PQM}$ for the default version. We find that the results are robust to the choice of the number of reference points and the distance metric. Any observed differences are within the standard deviation of the experiment.
@@
In this section, we study the effect of varying the two hyperparameters of our experiment: The number of reference points $n_{R}$, and the distance metric (which in all experiments in the main text is L2). We repeat the experiment on CIFAR10 described in § 3.3, varying the number of reference points, as well as changing the distance metric to L1. We also show the results when using the Inception-v3 feature extractor (Szegedy et al., 2016), as done in previous work such as Jiralerspong et al. (2023). We find that the order of the models is robust to the choice of the number of reference points and the distance metric, as well as to the use of a feature extractor. Any observed differences are within the standard deviation of the experiment. We show the results in Table 3.
@@ Appendix C Samples for validation
Figure 10: Samples from our the various sampling methods introduced in § 3.2, and their corresponding PQMass values.
In § 3.2, we used PQMass to compare samples from various sampling algorithms to true distribution samples. Fig. 10 shows the samples from each algorithm and the underlying distribution. The two distributions (gaussian mixture model and funnel) are often used for benchmarking sampling algorithms because of the high multimodality of the former and the complex shape of the latter. For the Gaussian mixture model (top), we see that MCMC is missing a mode, while the FAB samples are too noisy; for Neal’s funnel (bottom), we see that FAB’s samples look indistinguishable from true samples, while GGNS look similar but more spread out, while MCMC fails to model this distribution. In both cases, the results from $\chi^{2}_{\rm PQM}$ correlate well with the similarity to true samples we can see by eye.
@@ E.1 MNIST training progress
In the first part of § 3.3, we trained four generative models, two (identical) VAEs, and two (identical) diffusion models on the MNIST dataset for 100 epochs. In Fig. 5, we show the evolution of the $\chi^{2}_{\rm PQM}$ values as a training progresses. Here, in Fig. 11, we display one sample from run 1 of each generative model as a function of the training epoch to provide a qualitative understanding of the sensitivity of PQMass.
@@ E.2 Adding noise to ImageNet
In § 3.3, we evaluate PQMass’s and FLD’s ability to detect the addition of Gaussian noise with increasing variance compared to the test set from the low to high noise regime. As shown in Table 2, PQMass is more sensitive to noise corruption than FLD. We display an example showcasing the impact of each noise level on the sample in Fig. 12 to offer an intuitive understanding of the sensitivity of each metric.
@@ F.1 Gaussian mixture model
To evaluate our method’s ability to detect the diversity of generated samples, we reuse the Gaussian Mixture Model from § 3.1, but this time we systematically remove samples from $N$ modes, for varying $N$. For each $N$, we generate 5,000 samples and run our test 20 times with $n_{R}=100$, using different random reference points in each trial. The results are shown in Fig. 13. As expected, the $\chi^{2}_{\rm PQM}$ value increases as more modes are dropped, demonstrating that PQMass can effectively detect when a generative model fails to capture all modes, indicating reduced sample diversity.
@@ F.2 Detecting a small mode
In Table 4, we see that, as we include images from the new class, which introduces a small mode, PQMass can detect the difference between $X$ and $Y$ even for small $\alpha$. On the other hand, FLD struggles to detect the difference; FID also detects the difference (as shown by the monotonically increasing test statistic), but, unlike PQMass, does not give a point of reference allowing to compute significances (p-values).
@@ F.3 Time series
where $A$ is the amplitude of the signal and $\eta(t)$ is i.i.d. Gaussian noise with zero mean and unit variance. If $A=0$, then the time series is just noise. If $A\neq 0$, then there is a signal hiding in the noise. For each observation, we generate $100$ data points between $t=0$ and $t=10$. We then repeat the following process $5000$ times: We generate a time series with $A=0$ and a time series with $A\neq 0$. We then compare the two time series using PQMass, with $n_{R}=100$. We show the results in Fig. 14 for varying values of $A$. We can see that the $\chi^{2}_{\rm PQM}$ value increases as $A$ grows, as expected. The plot also the values of $\chi^{2}_{\rm PQM}$ corresponding to the $3\sigma$ and $5\sigma$ significance levels of detection. We see that, for this experiment, we can detect the signal with $5\sigma$ significance for $A\approx 0.12$, a signal that is invisible to the naked eye. We show an example of a time series with this amplitude, compared to one without signal in Fig. 15.
@@ F.4 Tabular data experiment
To show the versatility of our method on different types of data, we applied PQMass to the generation of tabular data. We used CTGAN Xu et al. (2019) as our generative model, trained on data from the Adult Census Dataset  [^6] https://archive.ics.uci.edu/dataset/2/adult. We convert categorical entries into a one-hot encoding for distance metrics between entries and compute distances between the encodings. We show the value of $\chi^{2}_{\rm PQM}$ in Fig. 16. We see that $\chi^{2}_{\rm PQM}$ goes down as training progresses, until eventually plateauing, as expected.
@@ F.5 Astrophysics experiments
As an example of PQMass applied to a scientific problem, we test the method on a problem from astrophysics. We trained a score-based model (SBM) on 256x256 real galaxy images from the Probes dataset, which is a compendium of 3163 high-quality local late-type galaxies Stone & Courteau (2019). Since there has been increasing interest in the use of SBMs as high-dimensional priors for inference problems, the question of the accuracy (in the statistical sense) of the trained SBM is very important. Using the PQMass metric, we were able to establish the probability that the generated samples come from the same underlying distribution as the training samples. On 256x256 images, PQMass runs in less than one hour for 1000 evaluations on a single Nvidia A100 GPU (about 3s for every $\chi^{2}$ evaluation). When we compare a random split of the probes dataset with itself, the $\chi^{2}_{\rm PQM}$ values are distributed according to the expected $\chi^{2}$ distribution while we can reproduce a curve similar to the one presented in Fig. 5 as a function of training iteration. The results are shown in Fig. 17.
Next, another scientific application of interest is using SBMs to model noise likelihoods as described in Legin et al. (2023). In this case, the accuracy of the learned generative model is of paramount importance. Here, we used the case of pixel darks (images of pure noise) from the Near Infrared Imager and Slitless Spectrograph (NIRISS) aboard the James Webb Space Telescope (JWST) as the noise to be modeled. This instrument has Gaussian noise as well as multiple sources of non-Gaussian additive noise, including dead pixels, faint and saturating cosmic rays, bleeding, a zero bias, blotches, and $1/f$ noise. The lattermost source of noise is especially difficult to completely correct for using traditional methods (Albert et al., 2023), effectively reducing the usable information of scientific images, especially for low signal-to-noise ratios. As an additional experiment, we have trained a SBM on 4000 of the full set of 6960 real 2048x256 pixel darks. We show examples of the ground truth images and generated images in Fig. 18 as well as two PQMass comparisons, one between two sets of the ground truth and another between the generated images and the ground truth. While the $\chi^{2}_{\rm PQM}$ metric can clearly show that half of the real dark samples are in distribution with the other half (bottom left plot in  Fig. 18) using 10 regions, it can also detect that the generated samples are not in distribution with real dark samples, despite their visually evident resemblance. Upon further investigation, this can be confirmed by the realization that the correct placement of hot pixels has not yet been fully learned by the generative model.
@@
We showcase our results in Table 5, where we use 10 regions for all experiments. In the null test, PQMass is consistently able to detect that they come from the same underlying distribution; however, FLD struggles, returning negative values indicating a measure of duplicates in the two datasets, which we know to not be true. FID struggles with Probes and SKIRT TNG but can showcase that for the JWST example, it is in distribution. Furthermore, when comparing samples generated from the SBM to the validation test, PQMass can detect that the samples come from the same distribution for Probes and SKIRT TNG, which we know to be correct. It can also capture the out-of-distribution nature of the JWST samples, which is correct. However, the same cannot be said for FLD or FID. For Probes and SKIRT TNG, FLD and FID claim the samples are out of distribution, and for JWST, FLD claims that the samples are in distribution, whereas FID correctly claims that the samples are out of distribution. Here, we showcase that PQMass can detect if the two datasets are in or out of distribution consistently where other standard metrics in the field either fail or are highly inconsistent.
@@
Probes (Ground Truth vs Ground Truth) 3x256x256 1029 vs 1030 $9.42$ $-21.72$ $84.64$
@@
SKIRT TNG (Ground Truth vs Ground Truth) 3x64x64 1230 vs 1229 $9.83$ $-4.09$ $12.14$
@@
JWST (Ground Truth vs Ground Truth) 1x256x2048 2000 vs 6969 $7.88$ $-3.06$ $1.07$
@@ F.6 Experiment on Protein Sequences
As another experiment on a real dataset with a different structure, we showcase the ability of PQMass to detect differences in datasets of protein sequences. As our example, we use the ESM Metagenomic Atlas dataset (Lin et al., 2023) [^7] https://esmatlas.com. We perform two comparisons: In the first one, we randomly split the dataset into two, and repeat the null test shown in § 3.1 and § A. As our distance metric, we use the Levenshtein, or edit distance (i.e., the number of edits needed to turn one sequence into the other). The result is shown in the blue histogram of Fig. 19. This shows not only that our method can also work in this data format, but also that it can work with a different distance metric.
Secondly, we split the dataset into two: One with Protein post-translational modification (PTM) higher than $0.5$, and one with PTM lower than $0.5$. This is equivalent to splitting the dataset into one of high confidence protein sequences, and one of low confidence. We then repeat the test and show the result in the orange histogram of Fig. 19. We can clearly see that PQMass can detect that these two datasets come from different distributions. It is a non-trivial result that, by comparing edit distances in sequence space, we can detect that the high PTM and low PTM sequences belong to different distributions.

### v3

@@ 4 Limitations
PQMass was shown to perform well with large number of samples. The statistics, however, will be too noisy for practical use if this number is too limited. Similarly, the number of regions is important: while our ablation study (§˜B) shows that the results are robust to various choices, in the limits $n_{R}=1$ and $n_{R}\rightarrow N$ (where $N$ is the number of samples), sensitivity will be limited.
@@
Finally, PQMass is only as good as the distance metric used, as shown in §˜G. For some modalities (for example, natural language), it is nontrivial to select a meaningful distance metric between samples.
@@ 5 Conclusion
We have shown the performance of PQMass on a variety of synthetic tasks, as well as on comparing sampling methods, comparing generative models of images, and detecting hidden signals in time series. Given the versatility (further shown in the additional experiments in §˜F) and low computational cost of PQMass, it can serve as a valuable tool for evaluating the quality and performance of generative models and sampling methods.
@@ Appendix A Null tests
Figure 9: Null tests, similar to the one shown in Fig.˜2, but for two different distributions: Samples from MNIST (left) and a VAE trained on MNIST (right). The main result of our paper holds for samples from any distribution as long as we are comparing two sets of independent samples from the same distribution. Note that, on MNIST, the fit is not perfect due to the limited number of data points.
The main result of this paper, shown in §˜2, was shown in practice in §˜3.1 for a high-dimensional mixture of Gaussians. In this section, we show that it works in even more complex and high-dimensional distributions. First, we test it on the MNIST data: Using the MNIST observations themselves as samples from some underlying probability distribution. We put together the MNIST train and test data, leading to $70,000$ images. We then split them into two subsets of $35,000$ images each. To repeat the null test of §˜3.1, we need independent samples at every iteration (note that this is different from the error bars reported in the rest of the experiments, which just arise from repeating with random tessellations, but always the same samples). Therefore, for each iteration, we take $1,000$ images from each dataset, and perform the PQM test with $n_{R}=100$, leading to 35 data points. We plot their histogram in the left panel of Fig.˜9. While noisy, due to the very few data points, we see that the histogram does follow a chi-squared distribution with $n_{R}-1$ degrees of freedom, as expected.
For our second experiment, we use one of the MNIST generative models described in §˜3.3. In this case, we can generate as many samples as we want. Because we are comparing the generative model to itself, the similarity of the generated samples to the MNIST train or test data is irrelevant for this particular test.
In practical applications, the number of samples available from real data might be limited, and in such cases, the PQMass distribution can be approximated by reusing the samples and only retessellating, which also yields good results. In practice, we have observed that, for typical sample sizes, the match between the $\chi^{2}_{PQM}$ distribution obtained by retessellating without resampling and the theoretical $\chi^{2}$ distribution is very close and more than adequate for testing the null hypothesis (we have indicated in the captions of the figures and the tables the specific experiments where this was done). However, in the cases where reusing the samples between retessellation, the retessellations are not independent since the samples are shared between each test, which might affect the obtained distribution. We explore further the impact of this approximation in §˜I.
@@ Appendix B Ablation study
Table 3: Ablation study for CIFAR-10. We repeat the experiment of §˜3.3, varying the number of reference points, as well as changing the distance metric to L1. We show the results for the frequentist version of PQMass, sorted in order of decreasing $\chi^{2}_{\rm PQM}$ for the default version. We find that the results are robust to the choice of the number of reference points and the distance metric. Any observed differences are within the standard deviation of the experiment.
@@
In this section, we study the effect of varying the two hyperparameters of our experiment: The number of reference points $n_{R}$, and the distance metric (which in all experiments in the main text is L2). We repeat the experiment on CIFAR10 described in §˜3.3, varying the number of reference points, as well as changing the distance metric to L1. We also show the results when using the Inception-v3 feature extractor (Szegedy et al., 2016), as done in previous work such as Jiralerspong et al. (2023). We find that the order of the models is robust to the choice of the number of reference points and the distance metric, as well as to the use of a feature extractor. Any observed differences are within the standard deviation of the experiment. We show the results in Table˜3.
@@ Appendix C Samples for validation
Figure 10: Samples from our the various sampling methods introduced in §˜3.2, and their corresponding PQMass values.
In §˜3.2, we used PQMass to compare samples from various sampling algorithms to true distribution samples. Fig.˜10 shows the samples from each algorithm and the underlying distribution. The two distributions (gaussian mixture model and funnel) are often used for benchmarking sampling algorithms because of the high multimodality of the former and the complex shape of the latter. For the Gaussian mixture model (top), we see that MCMC is missing a mode, while the FAB samples are too noisy; for Neal’s funnel (bottom), we see that FAB’s samples look indistinguishable from true samples, while GGNS look similar but more spread out, while MCMC fails to model this distribution. In both cases, the results from $\chi^{2}_{\rm PQM}$ correlate well with the similarity to true samples we can see by eye.
@@ E.1 MNIST training progress
In the first part of §˜3.3, we trained four generative models, two (identical) VAEs, and two (identical) diffusion models on the MNIST dataset for 100 epochs. In Fig.˜5, we show the evolution of the $\chi^{2}_{\rm PQM}$ values as a training progresses. Here, in Fig.˜11, we display one sample from run 1 of each generative model as a function of the training epoch to provide a qualitative understanding of the sensitivity of PQMass.
@@ E.2 Adding noise to ImageNet
In §˜3.3, we evaluate PQMass’s and FLD’s ability to detect the addition of Gaussian noise with increasing variance compared to the test set from the low to high noise regime. As shown in Table˜2, PQMass is more sensitive to noise corruption than FLD. We display an example showcasing the impact of each noise level on the sample in Fig.˜12 to offer an intuitive understanding of the sensitivity of each metric.
@@ F.1 Gaussian mixture model
To evaluate our method’s ability to detect the diversity of generated samples, we reuse the Gaussian Mixture Model from §˜3.1, but this time we systematically remove samples from $N$ modes, for varying $N$. For each $N$, we generate 5,000 samples and run our test 20 times with $n_{R}=100$, using different random reference points in each trial. The results are shown in Fig.˜13. As expected, the $\chi^{2}_{\rm PQM}$ value increases as more modes are dropped, demonstrating that PQMass can effectively detect when a generative model fails to capture all modes, indicating reduced sample diversity.
@@ F.2 Detecting a small mode
In Table˜4, we see that, as we include images from the new class, which introduces a small mode, PQMass can detect the difference between $X$ and $Y$ even for small $\alpha$. On the other hand, FLD struggles to detect the difference; FID also detects the difference (as shown by the monotonically increasing test statistic), but, unlike PQMass, does not give a point of reference allowing to compute significances (p-values).
@@ F.3 Time series
where $A$ is the amplitude of the signal and $\eta(t)$ is i.i.d. Gaussian noise with zero mean and unit variance. If $A=0$, then the time series is just noise. If $A\neq 0$, then there is a signal hiding in the noise. For each observation, we generate $100$ data points between $t=0$ and $t=10$. We then repeat the following process $5000$ times: We generate a time series with $A=0$ and a time series with $A\neq 0$. We then compare the two time series using PQMass, with $n_{R}=100$. We show the results in Fig.˜14 for varying values of $A$. We can see that the $\chi^{2}_{\rm PQM}$ value increases as $A$ grows, as expected. The plot also the values of $\chi^{2}_{\rm PQM}$ corresponding to the $3\sigma$ and $5\sigma$ significance levels of detection. We see that, for this experiment, we can detect the signal with $5\sigma$ significance for $A\approx 0.12$, a signal that is invisible to the naked eye. We show an example of a time series with this amplitude, compared to one without signal in Fig.˜15.
@@ F.4 Tabular data experiment
To show the versatility of our method on different types of data, we applied PQMass to the generation of tabular data. We used CTGAN Xu et al. (2019) as our generative model, trained on data from the Adult Census Dataset  [^6] https://archive.ics.uci.edu/dataset/2/adult. We convert categorical entries into a one-hot encoding for distance metrics between entries and compute distances between the encodings. We show the value of $\chi^{2}_{\rm PQM}$ in Fig.˜16. We see that $\chi^{2}_{\rm PQM}$ goes down as training progresses, until eventually plateauing, as expected.
@@ F.5 Astrophysics experiments
As an example of PQMass applied to a scientific problem, we test the method on a problem from astrophysics. We trained a score-based model (SBM) on 256x256 real galaxy images from the Probes dataset, which is a compendium of 3163 high-quality local late-type galaxies Stone & Courteau (2019). Since there has been increasing interest in the use of SBMs as high-dimensional priors for inference problems, the question of the accuracy (in the statistical sense) of the trained SBM is very important. Using the PQMass metric, we were able to establish the probability that the generated samples come from the same underlying distribution as the training samples. On 256x256 images, PQMass runs in less than one hour for 1000 evaluations on a single Nvidia A100 GPU (about 3s for every $\chi^{2}$ evaluation). When we compare a random split of the probes dataset with itself, the $\chi^{2}_{\rm PQM}$ values are distributed according to the expected $\chi^{2}$ distribution while we can reproduce a curve similar to the one presented in Fig. 5 as a function of training iteration. The results are shown in Fig.˜17.
Next, another scientific application of interest is using SBMs to model noise likelihoods as described in Legin et al. (2023). In this case, the accuracy of the learned generative model is of paramount importance. Here, we used the case of pixel darks (images of pure noise) from the Near Infrared Imager and Slitless Spectrograph (NIRISS) aboard the James Webb Space Telescope (JWST) as the noise to be modeled. This instrument has Gaussian noise as well as multiple sources of non-Gaussian additive noise, including dead pixels, faint and saturating cosmic rays, bleeding, a zero bias, blotches, and $1/f$ noise. The lattermost source of noise is especially difficult to completely correct for using traditional methods (Albert et al., 2023), effectively reducing the usable information of scientific images, especially for low signal-to-noise ratios. As an additional experiment, we have trained a SBM on 4000 of the full set of 6960 real 2048x256 pixel darks. We show examples of the ground truth images and generated images in Fig.˜18 as well as two PQMass comparisons, one between two sets of the ground truth and another between the generated images and the ground truth. While the $\chi^{2}_{\rm PQM}$ metric can clearly show that half of the real dark samples are in distribution with the other half (bottom left plot in  Fig.˜18) using 10 regions, it can also detect that the generated samples are not in distribution with real dark samples, despite their visually evident resemblance. Upon further investigation, this can be confirmed by the realization that the correct placement of hot pixels has not yet been fully learned by the generative model.
@@
We showcase our results in Table˜5, where we use 10 regions for all experiments. In the null test, PQMass is consistently able to detect that they come from the same underlying distribution; however, FLD struggles, returning negative values indicating a measure of duplicates in the two datasets, which we know to not be true. FID struggles with Probes and SKIRT TNG but can showcase that for the JWST example, it is in distribution. Furthermore, when comparing samples generated from the SBM to the validation test, PQMass can detect that the samples come from the same distribution for Probes and SKIRT TNG, which we know to be correct. It can also capture the out-of-distribution nature of the JWST samples, which is correct. However, the same cannot be said for FLD or FID. For Probes and SKIRT TNG, FLD and FID claim the samples are out of distribution, and for JWST, FLD claims that the samples are in distribution, whereas FID correctly claims that the samples are out of distribution. Here, we showcase that PQMass can detect if the two datasets are in or out of distribution consistently where other standard metrics in the field either fail or are highly inconsistent.
@@
Probes (Ground Truth vs Ground Truth) 3x256x256 1029 vs 1030 $9.42$ 21.72-21.72 $84.64$
@@
SKIRT TNG (Ground Truth vs Ground Truth) 3x64x64 1230 vs 1229 $9.83$ 4.09-4.09 $12.14$
@@
JWST (Ground Truth vs Ground Truth) 1x256x2048 2000 vs 6969 $7.88$ 3.06-3.06 $1.07$
@@ F.6 Experiment on Protein Sequences
As another experiment on a real dataset with a different structure, we showcase the ability of PQMass to detect differences in datasets of protein sequences. As our example, we use the ESM Metagenomic Atlas dataset (Lin et al., 2023) [^7] https://esmatlas.com. We perform two comparisons: In the first one, we randomly split the dataset into two, and repeat the null test shown in §˜3.1 and §˜A. As our distance metric, we use the Levenshtein, or edit distance (i.e., the number of edits needed to turn one sequence into the other). The result is shown in the blue histogram of Fig.˜19. This shows not only that our method can also work in this data format, but also that it can work with a different distance metric.
Secondly, we split the dataset into two: One with Protein post-translational modification (PTM) higher than $0.5$, and one with PTM lower than $0.5$. This is equivalent to splitting the dataset into one of high confidence protein sequences, and one of low confidence. We then repeat the test and show the result in the orange histogram of Fig.˜19. We can clearly see that PQMass can detect that these two datasets come from different distributions. It is a non-trivial result that, by comparing edit distances in sequence space, we can detect that the high PTM and low PTM sequences belong to different distributions.